OpenStudy (anonymous):

A rectangular sold has a volume of 640 cubic units. The width is 3 units more than the height and the length is 1 unit more than three times the height. Find the dimension

OpenStudy (zpupster):

why did you repost

OpenStudy (anonymous):

I didnt understand you

OpenStudy (zpupster):

i am sorry i hope someone can explain it better

OpenStudy (anonymous):

let height=p then length=3p+1 width=p+3 now it is given that (p)(3p+1)(p+3)=640 now solve the eqn and get the dimensions

OpenStudy (anonymous):

3p^3+10h^2+3p

OpenStudy (anonymous):

@ksrakhilesh now what?

OpenStudy (anonymous):

that is (p)(3p+1)(p+3)=5*16*8 this implies p=5

OpenStudy (anonymous):

how come you come up with that 5*16*8

OpenStudy (anonymous):

it is a mathematical technique to sole such equations by rearranging the right part untill it converges with left one

OpenStudy (anonymous):

how did you get p=5?

OpenStudy (anonymous):

@ksrakhilesh how did you get p=5?

OpenStudy (anonymous):

compare terms on both sides

OpenStudy (anonymous):